Ten yrs ago, a father was three times as old as his son. If he is only twice as old as his son, find their present ages?
Hope u help me answer this question? Tnx.

Answer by macston(5194)

F=father's age; S=son's age
F=2S
F-10=3(S-10) Substitute for F.
2S-10=3(S-10)
2S-10=3S-30 Add 30 to each side.
2S+20=3S Subtract 2S from each side.
20=S ANSWER 1: The son is 20 years old
F=2S=40 ANSWER 2: The father is 40 years old
CHECK
10 yrs ago, father was 3 times son's age:
F-10 yrs=3(S-10 yrs)
40 yrs-10yrs=3(20 yrs-10 yrs)
30 yrs=3(10 yrs)
30 yrs=30 yrs